IP One of the most popular events at Highland games is the hay toss, where competitors use a pitchfork to throw a bale of hay over a raised bar. Suppose the initial velocity of a bale of hay is v⃗ =(1.30m/s)x^+(9.35m/s)y^.After what minimum time is its speed equal to 4.00 m/s ?How long after the hay is tossed is it moving in a direction that is 45.0 ∘ below the horizontal?
Answer #1

Since horizonal velocity will remain same 'so

horizontal velocity= 1.3 m/s

for velocity to have magnitude 4 m/s ,let vertical velocity be v m/s

v^2+1.3^2=16

we got v=3.78 m/s

let the time is t

v=u-gt=>t=(u-v)/g=0.57 s

(b) it will be at 45 degree when both horizontal and vertical velocity are same

so vertical should be =-1.3 m/s

again using

v=u+gt

t=1.086 s